OpenStudy (anonymous):

There are 25 people competing in a race. In how many ways can they finish in first and second place? A)49 B)400 C)600 D)625

OpenStudy (black_ninja123):

There is 25 choices for first There is 24 choices for second Multiply the numbers and you should get your answer

OpenStudy (anonymous):

Thanks, could you help me with this? @Black_ninja123

OpenStudy (anonymous):

A bag contains tiles with the letters C-O-M-B-I-N-A-T-I-O-N-S. Lee chooses a tile without looking and doesn’t replace it. He chooses a second tile without looking. What is the probability that he will choose the letter O both times? A)1/132 B)1.72 C)1/66 D)1/23

OpenStudy (black_ninja123):

How many letters are there?

OpenStudy (black_ninja123):

How many o's are there?

OpenStudy (anonymous):

2 O's, 12 letter's

OpenStudy (anonymous):

@Black_ninja123

OpenStudy (black_ninja123):

\[\frac{ 2 }{ 12 }*\frac{ 1 }{ 11 }\]

OpenStudy (anonymous):

idk? @Black_ninja123

OpenStudy (black_ninja123):

you don't know how to multiply fractions?

OpenStudy (anonymous):

No....*blushes embarressed*

OpenStudy (black_ninja123):

\[\frac{ 1 }{ 6 }*\frac{ 1 }{ 11 }\] First Multiply the top numbers then multiply the bottom numbers

OpenStudy (anonymous):

1/66 is the answer?

OpenStudy (black_ninja123):

Yes

OpenStudy (anonymous):

ok, could you help me with 1 more question?

OpenStudy (anonymous):

The sample space for a roll of two number cubes is shown in the table. What is the probability that the roll will result in two odd numbers? A) 1/9 B) ¼ C) 1/3 D) 4/9

OpenStudy (black_ninja123):

How many sets have odd number in them?

OpenStudy (anonymous):

9?

OpenStudy (black_ninja123):

No

OpenStudy (anonymous):

?

OpenStudy (black_ninja123):

12

OpenStudy (black_ninja123):

wait one sec

OpenStudy (black_ninja123):

How many sets are there in total?

OpenStudy (anonymous):

36?

OpenStudy (black_ninja123):

so what is 12/36 simplified?

OpenStudy (anonymous):

1/3?

OpenStudy (black_ninja123):

Yes

OpenStudy (anonymous):

thanks!

OpenStudy (black_ninja123):

Your welcome :)